OpenStudy (anonymous):

Solve the following two equations simultaneously using substitution. 3x+2y=26 x-4y=4

OpenStudy (anonymous):

For substitution we need to make x = ... or y = ... lets take x-4y=4 Rearrange it so that it is in the form x = ...

OpenStudy (anonymous):

?

OpenStudy (anonymous):

(1)\[3x+2y=26\](2)\[x-4y=4\] Making (2) explicit for x:\[x=4y+4\]Substituting into (1):\[3(4y+4)+2y=26\]\[12y+12+2y=26\]\[14y=14\]\[y=1\]Substituting into (2):\[x-4(1)=4\]\[x=8\]